OpenStudy (anonymous):

Find the volume of a square pyramid with base 12 in by 12 in , height of 7.8 in, and slant height of 11 in. A)528 B)561.6 C)374.4 D)343.2

OpenStudy (anonymous):

7.8 * 12*12 ----------- 3

OpenStudy (lalaly):

Volume of pyramid =1/3 *Area of the base * Height area of base is area of square .... \[\text{volume of pyramid }=\frac{1}{3} \times 12 \times 12 \times 7.8\]

OpenStudy (anonymous):

C